Question 11

How much work is required to turn an electric dipole 180o in a uniform electric field of magnitude 41.0 N/C if p = 2.44 × 10-25 C·m and the initial angle is 59.3o.
Answer #1

The expression for the work done to rotate the dipole from the angle theta 1 to theta 2 is ,

W=-pE(cos(thetao+pi)-cos thetao)

=2pE cos thetao

=2(2.44*10^-25 C.m)(41 N/C) cos(59.3°)

=102.149*10^-25 J

=1.02*10^-23 J

Hence, required work is W=1.02*10^-23 J

I hope help you !!